Suppose that in the population, exactly 58% of all adults make a wish when they blow out candles on a birthday cake. If we select many random samples of size n = 100 people from this population, we know the sample proportion who make a wish when blowing out candles on a birthday cake will vary from sample to sample. The sampling distribution of the sample proportion would be Normal in shape and would have a standard deviation equal to what value?

A 0.1000

B 0.0494

C 0.0762

D 0.0024

E 0.5800
Answer #1

Ans:- P 581. 58/100 0,58 - I-P 10158 0:42 n = 100 - The standard deviation of Sample proportion (8= P(1-P) 615940142. loo 0.2